Sorry, but i need help with this one also. Thank you in advance.
Joe has a collection of nickels and dimes that is worth $5.65. If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45. How many dimes does he have?

Answer by stanbon(75887)

Joe has a collection of nickels and dimes that is worth $5.65.
5n + 10d = 565 cents
-------------------------------
If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45.
5(n+8) + 10(2d) = 1045 cents
5n+40 + 20d = 1045
-------------------------------------
How many dimes does he have?
-------------------------------------
EQUATIONS:
5n + 10d = 565
5n + 20d = 1005
----------------------
10d = 440
d = 44 (# of dimes)
-----------
Substitute into 5n+10d = 565 to solve for "n":
5n +440 = 565
5n = 125
n = 25 (# of nickels)
============================
Cheers,
Stan H.